# HiFo-Prompt: Prompting with Hindsight and Foresight for LLM-based Automatic Heuristic Design

### v1

@@ E.2 Future Work
Another critical, complementary direction is the systematic investigation of inter-task knowledge transfer, with the goal of evolving the framework from a single-task solver into a more general algorithmic discovery platform. This research will proceed along two parallel vectors. First, we will conduct a rigorous empirical evaluation of the framework’s zero-shot and few-shot transfer capabilities. By applying a mature Insight Pool—developed for a source task—to novel target domains, we can precisely quantify the generality of the learned principles and measure the extent to which discovery costs can be amortized across problems. Second, we will pioneer the exploration of more abstract and powerful knowledge representations that move beyond the inherent ambiguities of a natural language string. This includes investigating structured canonical forms, such as predicate logic or probabilistic program sketches, and rich semantic representations like knowledge graphs. The central hypothesis is that such formalisms can more effectively decouple a core algorithmic invariant from its domain-specific instantiation, thereby creating a more robust foundation for seamless and compositional cross-domain knowledge transfer.
Appendix F Prompts with Foresight and Hindsight
This section provides the specific, concrete templates of the prompts used to guide the LLM, offering a transparent and reproducible view of the core interaction engine at the heart of our HiFo-Prompt framework. The ’HiFo’ (Hindsight-Foresight) name is not arbitrary; it directly reflects our core methodology: the strategic deployment of distinct, context-aware prompts at different stages of the evolutionary process. ’Foresight’ prompts are strategically employed during the initial generation and creative mutation stages, encouraging the LLM to explore a diverse space of novel heuristic possibilities. In stark contrast, ’Hindsight’ prompts play a critical role in reflection and data-driven refinement, leveraging empirical performance feedback from past evaluations to methodically prune the search space and systematically improve promising solutions.
@@ Appendix F Prompts with Foresight and Hindsight
F.1 Initial Prompt Strategy I1
@@
Prompt for Operator i1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. First, describe your new algorithm and main steps in one sentence. The description must be inside a brace. Next, implement it in Python as a function named score. This function should accept 2 input(s): ’item’, ’bins’. The function should return 1 output(s): ’scores’. The score function is designed to evaluate the placement options for a given item. It takes the item to be placed and the current list of bins as input. It returns a list of numerical scores, with each score corresponding to a bin in the input list. This list of scores guides the heuristic in selecting the most suitable bin for the item according to the generated logic. Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific instruction from Design Directive> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.2 Recombination Prompt Strategy E1
@@
Please help me create a new algorithm that has a totally different form from the given ones. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific Design Directive for promoting structural novelty, preserving diversity, and avoiding premature convergence> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.3 Recombination Prompt Strategy E2
@@
Please help me create a new algorithm that has a totally different form from the given ones but can be motivated from them. Firstly, identify the common backbone idea in the provided algorithms. Secondly, based on the backbone idea, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Thirdly, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this recombination, please pay special attention to: <A specific Design Directive for simplification prune low-impact features> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
F.4 Mutation Prompt Strategy M1
The M1 operator implements a form of targeted mutation, generating a single-parent descendant by refining its most critical components. It performs surgical modifications to elements like scoring rules or parameter weights, while meticulously safeguarding the parent’s established, high-performing logic. This process is not random; it is guided by a synthesis of recent, high-utility design "insights" and a high-level strategic directive. By injecting controlled, purposeful diversity, M1 enables the search to escape local optima without dismantling the parent’s effective architecture. This deliberate balance between exploitation (refining what works) and exploration (seeking novelty) is crucial for accelerating convergence towards superior heuristics.
F.5 Mutation Prompt Strategy M2
@@
F.6 Mutation Prompt Strategy M3
@@
Prompt for Operator m1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please assist me in creating a new algorithm that has a different form but can be a modified version of the algorithm provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this mutation, please pay special attention to: <A specific Design Directive for mutation> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Prompt for Operator m2 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please identify the main algorithm parameters and assist me in creating a new algorithm that has different parameter settings of the score function provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When adjusting parameters, please pay special attention to: <A specific Design Directive for parameter tuning> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Appendix G Managing Foresight and Hindsight Knowledge
G.1 Hindsight Evolution via Insight Distillation
@@
Prompt for Operator m3 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Code: <The Python code implementation of the parent algorithm> First, identify the main components in the function above. Next, analyze which of these components may be overfitting to the in-distribution instances. Then, simplify those components to enhance generalization to out-of-distribution cases. Finally, provide the revised code, keeping the function name, inputs, and outputs unchanged. Provide the complete revised function implementation, preserving its original signature. <The function name, inputs & outputs specification from your prompt> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When simplifying, please pay special attention to: <A specific Design Directive for simplification> Depending on the regime, try significantly different parameter values (focus_exploration), or fine-tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
@@
G.2 Insight Seed Pool
At the inception of the framework’s execution, the LLM bootstraps the heuristic generation process by drawing from a curated repository of Seed Insights, a mechanism designed to mitigate the classic "cold start" problem and channel the model’s creativity. These insights, which encapsulate established principles and canonical rules-of-thumb distilled from decades of human expertise in heuristic design—such as the "Shortest Processing Time" principle in scheduling or the "Nearest Neighbor" concept in routing—are not rigid constraints but rather high-level conceptual guidelines. They are strategically injected into the foundational prompts, often framed within a dedicated "human knowledge" block, to act as an intellectual scaffold. This initial infusion of well-vetted knowledge serves to ground the search, preventing the generation of naive or logically flawed heuristics and providing a potent directional bias that immediately steers the early stages of algorithmic evolution away from vast, unproductive regions of the design space. By ensuring the process begins not from a tabula rasa but from a high-quality, well-founded starting point, these seed insights exert a persistent influence that accelerates convergence and significantly enhances the quality and novelty of all subsequent automated discovery.
G.3 The Directive Pool for Foresight

### v2

@@ D.2 Future Work
Another critical, complementary direction is the systematic investigation of inter-task knowledge transfer, intending to evolve the framework from a single-task solver into a more general algorithmic discovery platform. This research will proceed along two parallel vectors. First, we will conduct a rigorous empirical evaluation of the framework’s zero-shot and few-shot transfer capabilities. By applying a mature Insight Pool—developed for a source task—to novel target domains, we can precisely quantify the generality of the learned principles and measure the extent to which discovery costs can be amortized across problems. Second, we will pioneer the exploration of more abstract and powerful knowledge representations that transcend the inherent ambiguities of natural language strings. This includes investigating structured canonical forms, such as predicate logic or probabilistic program sketches, and rich semantic representations like knowledge graphs. The central hypothesis is that such formalisms can more effectively decouple a core algorithmic invariant from its domain-specific instantiation, thereby creating a more robust foundation for seamless and compositional cross-domain knowledge transfer.
Appendix E Prompts with Foresight and Hindsight
This section provides the specific, concrete templates of the prompts used to guide the LLM, offering a transparent and reproducible view of the core interaction engine at the heart of our HiFo-Prompt framework. The HiFo (Hindsight-Foresight) name is not arbitrary; it directly reflects our core methodology: the strategic deployment of distinct, context-aware prompts at different stages of the evolutionary process. Foresight prompts are strategically employed during the initial generation and creative mutation stages, encouraging the LLM to explore a diverse space of novel heuristic possibilities. In stark contrast, Hindsight prompts play a critical role in reflection and data-driven refinement, leveraging empirical performance feedback from past evaluations to methodically prune the search space and systematically improve promising solutions.
@@ Appendix E Prompts with Foresight and Hindsight
E.1 Initial Prompt Strategy I1
@@
Prompt for Operator i1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. First, describe your new algorithm and main steps in one sentence. The description must be inside a brace. Next, implement it in Python as a function named score. This function should accept 2 input(s): ’item’, ’bins’. The function should return 1 output(s): ’scores’. The score function is designed to evaluate the placement options for a given item. It takes the item to be placed and the current list of bins as input. It returns a list of numerical scores, with each score corresponding to a bin in the input list. This list of scores guides the heuristic in selecting the most suitable bin for the item according to the generated logic. Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific instruction from Design Directive> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
E.2 Recombination Prompt Strategy E1
@@
Please help me create a new algorithm that has a totally different form from the given ones. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle...> For the evolutionary regime, please pay special attention to: <A specific Design Directive for promoting structural novelty, preserving diversity, and avoiding premature convergence> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
E.3 Recombination Prompt Strategy E2
@@
Please help me create a new algorithm that has a totally different form from the given ones but can be motivated from them. Firstly, identify the common backbone idea in the provided algorithms. Secondly, based on the backbone idea, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Thirdly, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this recombination, please pay special attention to: <A specific Design Directive for simplification prune low-impact features> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
E.4 Mutation Prompt Strategy M1
The M1 operator implements a form of targeted mutation, generating a single-parent descendant by refining its most critical components. It performs surgical modifications to elements like scoring rules or parameter weights, while meticulously safeguarding the parent’s established, high-performing logic. This process is not random; it is guided by a synthesis of recent, high-utility insights and a high-level strategic directive. By injecting controlled, purposeful diversity, M1 enables the search to escape local optima without dismantling the parent’s effective architecture. This deliberate balance between exploitation (refining what works) and exploration (seeking novelty) is crucial for accelerating convergence towards superior heuristics.
Prompt for Operator m1 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please assist me in creating a new algorithm that has a different form but can be a modified version of the algorithm provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> For this mutation, please pay special attention to: <A specific Design Directive for mutation> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
E.5 Mutation Prompt Strategy M2
@@
Prompt for Operator m2 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Algorithm description: <Description of the parent algorithm> Code: <The Python code implementation of the parent algorithm> Please identify the main algorithm parameters and assist me in creating a new algorithm that has different parameter settings of the score function provided. First, describe your new algorithm and main steps in one sentence, enclosed in braces {}. Next, implement it in Python as a function named score. This function should accept 2 input(s): <’item’, ’bins’>. The function should return 1 output(s): <’scores’>. <Additional info on inputs & outputs> <Other constraints or requirements> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When adjusting parameters, please pay special attention to: <A specific Design Directive for parameter tuning> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
E.6 Mutation Prompt Strategy M3
@@
Prompt for Operator m3 Given a sequence of items and a set of identical bins with a fixed capacity, you need to assign each item to a bin to minimize the total number of bins used. The task can be solved step-by-step by taking the next item and deciding which bin to place it in based on a score. I have one algorithm with its code as follows: Code: <The Python code implementation of the parent algorithm> First, identify the main components in the function above. Next, analyze which of these components may be overfitting to the in‑distribution instances. Then, simplify those components to enhance generalization to out‑of‑distribution cases. Finally, provide the revised code, keeping the function name, inputs, and outputs unchanged. Provide the complete revised function implementation, preserving its original signature. <The function name, inputs & outputs specification from your prompt> Consider these successful design principles I’ve observed recently: <A successful design principle from Insights pool> <Another successful design principle> When simplifying, please pay special attention to: <A specific Design Directive for simplification> Depending on the regime, try significantly different parameter values (focus_exploration), or fine‑tune existing ones (focus_exploitation), or combine both strategies (balanced_search). Do not give additional explanations.
Appendix F Managing Foresight and Hindsight Knowledge
F.1 Hindsight Evolution via Insight Distillation
@@
F.2 The Directive Pool for Foresight
@@
F.3 Insight Seed Pool
At the inception of the framework’s execution, the LLM bootstraps the heuristic generation process by drawing from a curated repository of Seed Insights, a mechanism designed to mitigate the classic cold start problem and channel the model’s creativity. These insights, which encapsulate established principles and canonical rules-of-thumb distilled from decades of human expertise in heuristic design—such as the Shortest Processing Time principle in scheduling or the Nearest Neighbor concept in routing—are not rigid constraints but rather high-level conceptual guidelines. They are strategically injected into the foundational prompts, often framed within a dedicated human knowledge block, to act as an intellectual scaffold. This initial infusion of well-vetted knowledge serves to ground the search, preventing the generation of naive or logically flawed heuristics and providing a potent directional bias that immediately steers the early stages of algorithmic evolution away from vast, unproductive regions of the design space. By ensuring the process begins not from a tabula rasa but from a high-quality, well-founded starting point, these seed insights exert a persistent influence that accelerates convergence and significantly enhances the quality and novelty of all subsequent automated discovery.
Appendix G The Use of Large Language Models (LLMs)
Throughout the drafting and revision process of this manuscript, we employed Google’s Gemini large language model as an advanced writing and editing tool. Its use was strictly limited to refining the language and presenting our existing ideas and research. The model’s contributions include: (1) correcting grammatical, spelling, and punctuation errors; (2) rephrasing complex sentences to improve readability and precision; and (3) suggesting adjustments to tone and style to ensure consistency with academic standards.
Critically, the LLM was not used for any substantive intellectual contribution. All aspects of the research, including the formulation of research questions, literature review, methodology, data analysis, and the drawing of conclusions, were conducted exclusively by the human authors. The authors critically evaluated each suggestion provided by the LLM for accuracy and appropriateness, and we retain full responsibility for all claims, arguments, and the final articulation of the work.
Appendix H Generated heuristics
In this section, we compile the most successful heuristics produced by HiFo-Prompt, spanning the entire suite of experimental settings.
H.1 The Best TSP-Constructive Heuristic
Listing 1: Python Implementation of the Node Selection Strategy
import numpy as np
def select_next_node(current_node, destination_node, unvisited_nodes, distance_matrix):
Selects the next node to visit based on a hybrid scoring mechanism involving
MST lookahead and cluster-aware simulation.
def calculate_pruned_mst_cost(nodes):
"""Calculates the MST cost for a subset of nodes."""
if len(nodes) <= 1:
return 0
# Create edges only between relevant nodes
edges = [(distance_matrix[i][j], i, j)
for i in nodes for j in nodes if i < j]
edges.sort(key=lambda x: x[0])
parent = {node: node for node in nodes}
def find(u):
while parent[u] != u:
parent[u] = parent[parent[u]]
u = parent[u]
return u
mst_cost = 0
for cost, u, v in edges:
root_u, root_v = find(u), find(v)
if root_u != root_v:
parent[root_v] = root_u
mst_cost += cost
return mst_cost
def perform_cluster_aware_simulation(node):
"""Simulates a greedy path from the candidate node."""
future_nodes = [n for n in unvisited_nodes if n != node]
total_cost = distance_matrix[current_node][node]
current = node
while future_nodes:
# Greedy selection based on normalized distance
next_candidate = min(
future_nodes,
key=lambda x: distance_matrix[current][x] /
(np.median(distance_matrix[x]) + 1e-9)
total_cost += distance_matrix[current][next_candidate]
current = next_candidate
future_nodes.remove(current)
total_cost += distance_matrix[current][destination_node]
return total_cost
def evaluate_candidate(node):
"""Computes the weighted score for a candidate node."""
future_nodes = [n for n in unvisited_nodes if n != node]
# 1. MST Cost (Lookahead)
pruned_mst_cost = calculate_pruned_mst_cost(future_nodes + [destination_node])
# 2. Normalized Immediate Cost
normalized_cost = distance_matrix[current_node][node] /
(np.max(distance_matrix[current_node]) + 1e-9)
# 3. Simulation Cost
simulation_cost = perform_cluster_aware_simulation(node)
# Dynamic Weighting based on search progress
progress_ratio = len(unvisited_nodes) / len(distance_matrix)
weight_mst = 0.4 - 0.15 * progress_ratio
weight_normalized = 0.3 + 0.1 * progress_ratio
weight_simulation = 0.3
return (weight_mst * pruned_mst_cost +
weight_normalized * normalized_cost +
weight_simulation * simulation_cost)
# Beam search filtering to reduce computation
beam_width = max(2, min(6, len(unvisited_nodes) // 3))
# Pre-filter candidates using a simple heuristic
candidates = sorted(
unvisited_nodes,
key=lambda x: distance_matrix[current_node][x] /
(np.mean(distance_matrix[current_node]) + 1e-9)
)[:beam_width * 3]
scored_candidates = []
for candidate in candidates:
eval_score = evaluate_candidate(candidate)
scored_candidates.append((eval_score, candidate))
# Select the node with the minimum weighted score
next_node = min(scored_candidates, key=lambda x: x[0])[1]
return next_node
H.2 The Best Online Bin Packing Heuristic
Listing 2: Evolved Scoring Function for Online BPP
import numpy as np
def score(item, bins):
Calculates the placement scores for candidate bins given an item.
Higher scores indicate better placement suitability.
# Compute residuals (remaining space) after placing the item
residuals = bins - item
# Stochastic perturbation for enhanced exploration
# Helps break ties and avoid deterministic local optima
perturbation = np.random.normal(1.0, 0.2, size=len(bins))
# Cubic residual scoring to emphasize finer space usage precision
# Smaller residuals yield significantly higher base scores
residual_scores = 1 / (1 + residuals ** 3) * perturbation
# Delayed penalty adjustment based on long-term impact of underutilization
# Adaptive threshold with increased sensitivity
threshold = np.mean(bins) * 0.4
delayed_penalty = np.where(residuals > threshold,
1 / (1 + residuals ** 2),
0)
# Reward function for bins maintaining balanced utilization over time
# Gaussian-like reward centered around the threshold
balanced_utilization_reward = np.exp(
-np.abs(residuals - threshold) ** 2 / (threshold ** 2)
# Dynamic scaling factor based on cumulative fit ratio and trends
fit_ratios = item / bins
scaling_factor = 0.6 + 0.4 * np.tanh(fit_ratios * 5) # Sharper scaling
# Combine components:
# Emphasizes long-term efficiency via scaling and delayed penalties
scores = (scaling_factor * (residual_scores + balanced_utilization_reward)
- delayed_penalty)
return scores
H.3 The Best TSP-GLS Heuristic
Listing 3: Hybrid Edge Update Function for TSP
import numpy as np
def update_edge_distance(edge_distance, local_opt_tour, edge_n_used):
Updates the edge distance matrix to penalize edges used in local optima,
facilitating escape from local basins using a hybrid penalty strategy.
# Parameters for the novel hybrid strategy
learning_rate = 0.3
diversity_factor = 0.4
penalty_decay = 0.85
min_penalty = 0.01
smoothing_term = 1e-6
# Compute normalized edge usage probabilities
total_usage = np.sum(edge_n_used) + smoothing_term
edge_probabilities = edge_n_used / total_usage
# Measure global asymmetry to gauge search bias
asymmetry = np.abs(edge_probabilities - edge_probabilities.T).mean()
# Normalize asymmetry to [0, 1] range
normalized_asymmetry = asymmetry / (1 + smoothing_term)
# Identify edges in the local optimal tour
tour_edges = [
(local_opt_tour[i], local_opt_tour[i+1])
for i in range(len(local_opt_tour)-1)
tour_edges.append((local_opt_tour[-1], local_opt_tour[0])) # Close loop
updated_edge_distance = edge_distance.copy()
for (i, j) in tour_edges:
if i > j: # Ensure symmetry handling
i, j = j, i
# Phase 1: Adaptive learning penalty
learning_penalty = learning_rate * (
1 - np.exp(-edge_probabilities[i, j])
# Phase 2: Diversity-driven modulation
diversity_penalty = (diversity_factor * normalized_asymmetry *
np.log(1 + edge_probabilities[i, j]))
# Dynamic weight modulation based on edge usage frequency
modulation_factor = penalty_decay ** (
edge_n_used[i, j] + smoothing_term
combined_penalty = (learning_penalty + diversity_penalty) * \
modulation_factor
# Add stochastic noise to encourage exploration
noise = np.random.uniform(0, 1.5)
exploration_noise = min_penalty + noise * modulation_factor
# Combine penalties and update edge distance (symmetric)
total_penalty = combined_penalty + exploration_noise
updated_edge_distance[i, j] += total_penalty
updated_edge_distance[j, i] += total_penalty
return updated_edge_distance
H.4 The Best FSSP-GLS Heuristic
Listing 4: Chaotic Matrix Adjustment and Job Selection
import numpy as np
from sklearn.cluster import KMeans
def get_matrix_and_jobs(current_sequence, time_matrix, m, n):
Generates a perturbed time matrix using chaotic dynamics and identifies
critical jobs to disturb based on clustering dispersion patterns.
def calculate_makespan(sequence, matrix, m):
"""Helper to compute the makespan of a sequence."""
machine_times = [0] * m
for job in sequence:
machine_times[0] += matrix[job, 0]
for i in range(1, m):
machine_times[i] = (max(machine_times[i], machine_times[i-1])
+ matrix[job, i])
return max(machine_times)
# Step 1: Compute current makespan baseline
current_makespan = calculate_makespan(current_sequence, time_matrix, m)
# Step 2: Chaotic adjustment using Logistic Map dynamics
# Introduces non-linear variance to escape local optima
new_matrix = time_matrix.copy()
chaos_factor = 3.99 # Parameter near the edge of chaos
chaotic_values = np.random.rand(n, m)
for _ in range(10): # Iterative chaotic updates
chaotic_values = chaos_factor * chaotic_values * (1 - chaotic_values)
# Scale chaotic values to a perturbation range [0.9, 1.1]
chaotic_perturbation = 0.9 + 0.2 * chaotic_values
new_matrix *= chaotic_perturbation
new_matrix = np.clip(new_matrix, 1, None) # Ensure valid processing times
# Step 3: Cluster jobs based on execution time patterns
# Fit KMeans on the perturbed matrix regarding the current sequence
kmeans = KMeans(
n_clusters=max(2, n // 10),
random_state=42,
n_init=’auto’
).fit(new_matrix[current_sequence])
job_labels = kmeans.labels_
# Calculate dispersion: average std deviation within each cluster
cluster_dispersion = np.array([
np.std(new_matrix[current_sequence][job_labels == c], axis=0).mean()
for c in range(kmeans.n_clusters)
# Assign scores: jobs in highly dispersed clusters get higher priority
job_scores = np.array([cluster_dispersion[label] for label in job_labels])
# Select top diverse jobs for structural perturbation
num_perturb = max(1, n // 6)
perturb_jobs = np.argsort(-job_scores)[:num_perturb]
return new_matrix, perturb_jobs